### v4

# Sparrow: Data-Efficient Video-LLM with Text-to-Image Augmentation

Abstract
Recent years have witnessed the success of Multimodal Large Language Models (MLLMs) in the vision understanding domain. The success of these models can largely be attributed to the dominant scaling law, which states that larger parameter sizes and data volumes contribute to better performance. Notably, data scaling has mainly been powered by automatic data pipelines, which center around the self-instruction of LLMs. The paradigm has been taken for granted for quite some time, but the study of the effectiveness of scaling with these data has been neglected for a long time. In this context, this work revisits scaling with synthetic data and focuses on developing video-LLMs from a data-centric perspective. Our main study approach is fine-tuning pre-trained image-LLMs with video data and investigating learning efficiency through data scaling. Results from our preliminary experiments reveal a low learning efficiency phenomenon when simply scaling up video data samples, which, through our probing, can be ascribed to a lack of instruction diversity. Aiming at this issue, we propose a data augmentation method called Sparrow, which synthesizes video-like samples from pure text instruction data. Mixing these synthetic samples with the video data enables a more efficient training scheme. Through comprehensive experiments, we demonstrate that our proposed method achieves performance comparable to or even superior to baselines trained with many more samples. Meanwhile, we find that incorporating these synthetic samples can boost the performance of long video understanding without training with long video data. The code and data examples are available at this link.
Figure 1: Performance comparison with different schemes of data scaling. We introduce Sparrow, a data-efficient training method for video-LLMs that achieves “Train less, gain more” in both general video and long video understanding performance. 0 sample indicates zero-shot inference with the image-LLM [1].
1 Introduction
The past few years have seen the rapid progress of Multimodal Large Language Models (MLLMs) [2, 3, 4]. Apart from solving traditional vision tasks (such as VQA), these models also excel in following user instructions and generalizing to new tasks. A mainstream paradigm for developing such models takes a two-stage training strategy. The first stage, pretraining, mainly serves to align vision modality with text and inject various kinds of visual knowledge into the model. In this stage, large-scale datasets of text-image pairs are often used, such as LAION [5] and CC [6], comprising a large proportion of the total compute and injecting abundant vision knowledge into models. Some methods also incorporate OCR and detection-related data to improve foundational capabilities [7, 4]. The second stage, instruction fine-tuning, adapts models to accommodate various tasks and helps generalize to new instructions. Training in this stage typically involves instruction data obtained from self-instruction or adaptation of task-specific datasets (e.g. VQA and chart understanding datasets). Recently, researchers have shifted their focus from single-image models to more advanced ones that support video understanding. Borrowing successful experience from developing image models, some video counterparts are typically trained from scratch, following a similar two-stage training paradigm [8, 9]. Apart from this path, some researchers utilize pre-trained image-LLMs instead. Typical approaches include zero-shot inference [10, 11, 12] and further fine-tuning [13, 9, 14, 8].
Notably, the success of these models can be largely ascribed to the formidable scaling law, which puts emphasis on scaling up parameter size or data volume for better model performance. For the data aspect, the scaling has mainly been driven by automatic data engines, which synthesize massive amounts of data without human labor. Nevertheless, the characteristics of learning from these synthesized video data stand out as a critical yet underexplored topic. Thus, in this work, we investigate the learning characteristics of video-LLM more deeply from a data scaling perspective. Our preliminary data scaling experiments reveal a low data efficiency problem, that is, the performance gains from utilizing multiple times more data is marginal. An inspection of data characteristics suggests this might be due to a lack of instruction diversity in the training corpus. To address this issue, we propose a data augmentation method, dubbed Sparrow [^1] Inspiration taken from the swiftness of sparrows., to enrich the diversity of instruction. The basic idea is to synthesize video-like samples from textual data and mix these synthetic data with the video samples. Specifically, we use existing text instruction data whose sample comprises a (long-context, instruction, answer) triplet. The long-context part is split into multiple segments and then further transformed into images, while the instruction and answer stay intact. Processed in this way, the synthetic samples have the same structure as video instruction data and can be incorporated seamlessly into the training corpus.
Comprehensive experiments demonstrate that our methods can facilitate data-efficient fine-tuning of image-LLMs for general video understanding and assist models in the comprehension of long videos. Specifically, using the same number of training samples, our method shows clear advantages over other data schemes. It even surpasses the baselines trained with many more samples, achieving high data efficiency (Fig. 1). The contributions of this work include:
  • We investigate the fine-tuning approach for developing video-LLMs from a data perspective, and shed light on possible factors that lead to low learning efficiency.
  • We propose a data augmentation method that improves the instruction diversity of training data and facilitates a more efficient training scheme.
  • We perform comprehensive experiments to evaluate the proposed method and examine its key properties, paving the way for future research in this line.
2 Related Work
2.1 Multimodal Large Language Models
Image-LLMs. To develop image-LLMs, the mainstream approach is to build upon powerful pre-trained LLMs and extend LLMs with the capability to perceive and reason with images [3, 4]. Based on a two-stage training recipe, i.e. image-text alignment training and instruction tuning, the developed model can fulfill a wide range of multimodal user queries and present its answers in user-friendly natural language sentences.
Video-LLMs. Following the success of image-LLMs, subsequent endeavors aim to expand the triumph to more intricate video understanding. Works like Video-ChatGPT [15], VTimeLLM [16], PLLaVA [17] and LLaVA-NeXT-Video [18] attempt further fine-tune image-LLMs to enhance video understanding capability. Other research [13, 9, 14, 8] explores training from pre-trained LLM, following the basic alignment-then-finetuning paradigm similar to image-LLM. These approaches usually involve joint training that mixes image and video data in the training corpus. In this study, we build upon pre-trained image-LLMs and enhance video understanding capabilities through fine-tuning.
2.2 Evaluation of Video Understanding
Early methods [10, 15] are generally evaluated on more traditional benchmarks like MSVD-QA [19], TGIF-QA [20] and ActivityNet-QA [21]). These benchmarks are generally domain-specific and focus on certain basic skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage (especially in longer videos) and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning.
Recently, with the rise of benchmarks specifically designed for MLLMs [22, 9, 23, 24], a more in-depth and comprehensive evaluation has become more accessible. Compared to previous traditional benchmarks, these newly developed benchmarks are generally more challenging, often entailing composite skills and a finer-grained understanding of the video (e.g. the plot in the movie or causal relationships between events), and can be much longer in duration (e.g. up to 60 minutes in the Video-MME benchmark). In this work, our study adopts these newly developed video benchmarks.
2.3 Textual Data for Video Understanding
Since MLLMs are typically built upon LLMs and thus highly compatible with textual data, some works have explored utilizing pure text data to boost the performance of video understanding. Below, we outline the key ideas of these methods and their differences from our method.
  • Textual data for context expanding. Previous works have explored utilizing textual data to expand the context window of base LLMs. Specifically, in order to facilitate long video understanding, LLaMA-VID [25] and LongVA [26] incorporate long text data in fine-tuning and continued pre-training stages, respectively, to expand the context window of LLM backbones.
    Differences: In this work, we (1) adopt textual data as a data augmentation method and (2) use them in the vision form to accommodate the training format.
  • Synthetic textual data for video understanding. This line of work investigates synthesizing textual data that simulates video QA data, aiming to transfer temporal reasoning capabilities from textual training. More specifically, TOPA [27] extracts textual captions and object-level information from video frames, while T3 [28] gathers similar information from multiple different images.
    Differences: These two works seek to boost video understanding with synthetic data, while ours aims to enrich the instruction diversity of the training corpus. Moreover, our method does not require calling advanced LLM APIs to build data; instead, our method utilizes existing datasets.
3 A Probing Study of Data Scaling
To understand the scaling characteristics of training data, our study starts with fine-tuning with different sample sizes and examines the relationship between training sample size and model performance.
In this section, we introduce the study’s training and evaluation setup and then illustrate the empirical findings.
3.1 Training Setup
3.1.1 Model Setup
During our exploration, we mainly utilize two image-LLMs, including Mini-InternVL-Chat-4B-V1.5 [7] (termed as InternVL hereafter), MiniCPM-Llama3-8B-V2.5 [1] (termed as MiniCPM-8B hereafter). These instruction-tuned models are trained with massive image data and equipped with strong foundational capabilities. To support higher-resolution vision input, these models adopt the patchifying technique [29, 30, 31] with a dynamic resolution scheme, where an image can be cropped into multiple sub-images according to different aspect ratios. Specifically, InternVL supports up to 13 sub-images, each of which is converted into 256 visual tokens; MiniCPM-8B slices images into a maximum of 10 patches, where each is represented by 96 visual tokens. During training and evaluation, we switch off the patchifying option for higher efficiency.
3.1.2 Training Configurations
For fairness and ease of reproduction, we follow the official implementations. More specifically, we train the whole model end-to-end (except for InternVL-4B, where we freeze the vision encoder) with a learning rate of 5e-6.
(a) Share-Gemini
(b) Video-ChatGPT
Figure 2: Video length statistics of ShareGemini and Video-ChatGPT datasets. Both datasets mostly cover videos shorter than 3.5 minutes. We extract video frames at an FPS of 1 for each video.. For better visibility, we pick samples with frame numbers lower than 99.9 percentile for visualization.
3.1.3 Training Datasets
During our investigation, we utilize two representative types of datasets, i.e., video-caption pairs and video instruction data. Specifically, we choose the ShareGemini [32] dataset and the Video-ChatGPT [15] dataset as caption and instruction data, respectively. For each video, frames are extracted at an FPS of 1. In consideration of efficiency, we use up to 64 frames for InternVL-4B and 24 frames for MiniCPM-8B. When the total number of frames exceeds the threshold, we uniformly downsample the video frames. The statistics of video lengths are shown in Fig. 2, and we provide more introduction to the two datasets below.
ShareGemini-Webvid-core100k. It is a video caption dataset with 100K samples in total. The videos are curated from WebVid [33], a web-scale video-caption dataset covering open domains and general subjects. Regarding duration, the dataset mainly contains short videos with lengths shorter than 30 seconds.
The captions are annotated by calling the strong Gemini-1.5-Pro [34] API. To ensure the diversity of video content, an advanced clustering algorithm [35] is used to filter out highly similar videos. For simplicity, we refer to this dataset as ShareGemini in the following parts of the paper.
Video-ChatGPT. The video instruction dataset contains 100K video-instruction pairs. The videos in this collection are derived from ActivityNet [36]. The dataset’s coverage of video duration is larger, yet the average video length is no more than 3.5 minutes. There are broadly three types of instructions: video summarization, questions about video content, and creative/generative tasks.
The dataset is annotated in a semi-automatic manner. A small portion of data samples are manually annotated by human annotators by refining and enriching the video captions. Other instruction data are generated by GPT-3.5 with the aid of off-the-shelf dense prediction and captioning models.
3.2 Evaluation Setup
To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [22], MVBench [9], and TempCompass [23] as our benchmarks. We do not use traditional video-QA benchmarks (e.g. MSVD-QA [19], TGIF-QA [20], ActivityNet-QA [21]) since these benchmarks are generally limited to a small coverage of domains, task types, and video lengths. Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel. We illustrate more about the benchmarks used as follows:
Video-MME is a comprehensive benchmark designed for the evaluation of video-LLMs. For temporal coverage, videos of short length (up to 2 minutes), medium length (4–15 minutes), and longer duration (30–60 minutes) are included. The videos and annotations are manually collected and filtered. We only use the raw frames without the subtitles to focus on the evaluation of video understanding capabilities.
MVBench designs a set of 20 video tasks that cover both perception and cognition, such as scene transition and episodic reasoning. Compared to Video-MME, the videos are sourced from existing benchmarks, and the QAs are automatically generated for the 20 pre-defined tasks.
TempCompass focuses on fine-grained temporal aspects, such as action, speed, and attribute change. The videos and meta-information are manually collected, after which annotations are generated by LLMs with the aid of meta-information. We use the multiple-choice QA (MCQ) format to align with other benchmarks.
To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment. More details about the implementation of this evaluation scheme are available in Appendix A.1.
3.3 Main Findings
3.3.1 Low Learning Efficiency Issue
Figure 3: Scaling performance of fine-tuning InternVL with different data volumes and types on different benchmarks. We consider fine-tuning with ShareGemini, Video-ChatGPT, and a mix of both with a 1:$1:11$ : 1 sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference.
Figure 4: t-SNE plot of instruction distribution of video datasets - ShareGemini and Video-ChatGPT. 5,000 instructions are sampled from each dataset. The instruction diversity of these two datasets is relatively limited, leading to low learning efficiency.
Our experiments start with scaling up the training data volume and evaluating the video understanding performance on different general video understanding benchmarks. The results are shown in Fig. 3. In general, training either with video caption data (ShareGemini), instruction data (Video-ChatGPT), or a mix of both can boost the image-LLM’s video understanding performance. Meanwhile, increasing the training volume brings additional gains in accordance with the data scaling law. However, the gains from scaling up quickly reach a plateau. For instance, on the Video-MME benchmark, when training with mixed data, 30K samples improve overall accuracy by 3.1 points, while 100K samples only add another 0.5 points, resembling a logarithmic growth. In view of this quick and early saturation, the learning efficiency with these video datasets can be quite limited. The phenomenon also suggests that there could be high redundancy in the training corpus, and it is possible that we may use less data to achieve a performance comparable to or even better than training with more data samples.
3.3.2 Probing of Instruction Diversity
Previous results prompt us to explore the reason for such low learning efficiency. Inspired by prior studies, which have underscored the importance of instruction diversity for fine-tuning LLMs [37] and image-LLMs [38], we conduct an inspection of training data in this aspect. Specifically, we follow previous approaches [39, 40] to visualize the distribution of instructions in the training corpus. 5,000 instructions are sampled from ShareGemini and Video-ChatGPT, respectively. Then, the instructions are embedded and visualized using the t-SNE technique, as shown in Fig. 4. Overall, the instruction distribution of these two datasets is not diverse enough, which leads to a low data efficiency: The distribution of ShareGemini exhibits 9 clear clusters in the figure, indicating very similar instructions. This is because this dataset samples from a fixed pool of 9 templates as instructions, each of which is a variant of “Describe this video in detail”. On the other hand, the distribution of Video-ChatGPT seems relatively more diverse, as it includes specific questions related to video content and details besides video summarization. Nevertheless, the instruction diversity is still low due to the nature of self-instruction and a few fixed task-specific prompting templates for data curation.
4 Methods
4.1 Design Concept
Figure 5: A conceptual illustration of our scheme for synthesizing video-like samples from text data. For each (long-context, instruction, answer) triplet, we split the long context into segments and transcribe these segments into a series of images, simulating the structure of a video. Zoom in for better visibility.
Since currently available video data can be limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity by incorporating new synthetic data. A rich source of instruction data lies in the text domain, and they can effectively complement the vision domain. Nevertheless, there is inherently a modality gap between the text and visual domains. To better utilize these data, we bridge the modality gap by synthesizing images with the text. Fig. 5 illustrates our overall workflow for synthesis of data.
Figure 6: t-SNE plot of instruction distribution after applying our proposed methods. By incorporating text data, i.e. LongAlpaca [41] and LongQLoRA [41], the overall distribution becomes more diverse.
Our proposed scheme enjoys three benefits: (1) Mixing in text data can effectively enrich the instruction diversity (Fig. 6), thus improving the learning efficiency for video fine-tuning; (2) Images synthesized from text can emulate the 1D temporal structure of video frames since text segments are generally correlated in the context, thus mitigating the gap between common video samples and synthetic ones; (3) Text data are easier to collect than video samples. Thus, utilizing synthetic data can be economical.
4.2 Implementation Details
Each text sample is a (long-context, question, answer) triplet. For example, the long context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g. an inquiry to give a synopsis, or questions related to the paper. After the data transformation process, each sample is a video-like (images, question, answer) triplet, where long-context information is transformed into a series of images, and the question and answer stay unchanged.
The key to the data synthesis procedure is synthesizing images with pure text. Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments according to word counts (set to 115 empirically) using an open-sourced NLP toolkit [^2] https://www.nltk.org/. These text chunks are then transformed into a sequence of images separately. Specifically, each text chunk is embedded into a blank image with a white background. This is achieved using bitmap font with the ImageFont module of Pillow library [^3] https://pillow.readthedocs.io/en/stable/. Each image is 448x448 pixels in size, and the font is 20 pt large, black color, Arial Regular type. We use a bounding box to control the layout, leaving a margin of 20 pixels on each side, so each line of the text has roughly the same width. After the transformation, the structure of these synthesized data is exactly the same as that of video samples, and we can seamlessly incorporate the synthetic data into the video training corpus.
5 Evaluation on Proposed Methods
This section includes experimental results and discussions of our proposed method, including (1) a comparison with mainstream methods, (2) an ablation study on data mixes, and (3) an examination of key properties, including data scaling performance and gains in long video understanding.
Methods Size Frames Short Medium Long Overall
Proprietary Models
GPT-4V [42] N/A 10 70.5 55.8 53.5 59.9
Claude 3.5 Sonnet [43] N/A 20 71.0 57.4 51.2 60.0
GPT-4o [44] N/A 384 80.0 70.3 65.3 71.9
Gemini 1.5 Pro [45] N/A 1fps 81.7 74.3 67.4 75.0
Open-Source Models
VideoChat2 [9] 7B 16 48.3 37.0 33.2 39.5
Video-LLaVA [13] 7B 8 45.3 38.0 36.2 39.9
Chat-UniVi-v1.5 [14] 7B 64 45.7 40.3 35.8 40.6
VideoLLaMA 2 [8] 7B 16 56.0 45.4 42.1 47.9
VITA [46] 8x7B 32 65.9 52.9 48.6 55.8
Kangaroo [47] 8B 64 66.1 55.3 46.6 56.0
VITA-1.5 [48] 7B 16 67.0 54.2 47.1 56.1
FT w/ InternVL [7]
Zero-shot 3.8B 64 61.3 51.8 44.3 52.5
200K video data 3.8B 64 66.7 54.2 48.1 56.3
Sparrow (30K hybrid data) 3.8B 64 67.0 53.7 49.3 56.7
Table 1: Accuracy comparisons of different methods on the Video-MME benchmark. Performance is ranked in ascending order regarding overall performance (The models are grouped according to open-source or not). Our method uses only 15% of the total sample size compared to the full volume video datasets (200K) for fine-tuning and achieves comparable performance. 30K hybrid data comprise 20K data sampled from video datasets and 10K synthesized from our method. Bold digits indicate the best performance within each group.
5.1 Comparison with Mainstream Methods
We compare the results with some representative proprietary models, including GPT-4V [42], Claude 3.5 Sonnet [43], GPT-4o [44], Gemini 1.5 Pro [45], and open-source video-LLMs of similar LLM parameter size, including Video-LLaVA [13], VideoChat2 [9], Chat-UniVi-v1.5 [14], VideoLLaMA 2 [8], VITA [46], VITA-1.5 [48], and Kangaroo [47], as summarized in Tab. 1.
The table results show that, through zero-shot inference, the image-LLM Intern-VL already outperforms a variety of video-LLMs with larger LLM parameter sizes. This might be due to the rich pre-trained knowledge embedded in the model parameters since the image-LLM has been trained with large-scale and high-quality image-text data. The vision prior lays a strong foundation for further video fine-tuning, where models learn temporal and causal concepts from activities, events, etc. The model fine-tuned with full video datasets achieves an overall gain of 3.8 points on the image-LLM, closing the gap between open-source models and proprietary ones.
Notably, our methods use only 15% of the total sample size compared to the full volume (200K) for fine-tuning and achieve comparable performance. This result suggests the high data efficiency of our proposed scheme since mixing in synthetic samples mitigates the low instruction diversity issue illustrated in the earlier section.
5.2 Ablation on Different Data Compositions
Data Mix S M L Overall
30K Share-Gemini 65.7 52.8 46.1 54.9
30K Video-ChatGPT 66.3 53.0 47.3 55.6
15K Share-Gemini 15K Video-ChatGPT 66.2 53.3 47.4 55.7
10K Share-Gemini 10K Video-ChatGPT 10K synthetic 67.0 53.7 49.3 56.7
10K Share-Gemini 10K Video-ChatGPT 10K pure text 67.3 52.4 47.7 55.8
Zero-shot 61.3 51.8 44.3 52.5
200K full data 66.7 54.2 48.1 56.3
Table 2: Results of different data compositions on the Video-MME benchmark. Our proposed scheme achieves an overall performance superior to other data mixes of the same amount (30K) and even more data (200K). The Zero-shot and 200K full data fine-tuning settings are listed for comparison.
In order to examine the impact of different data compositions and validate the effectiveness of the proposed method, we conduct an ablation study and construct the following settings with the same amount of total data samples:
  • 30K video samples from ShareGemini.
  • 30K video samples from Video-ChatGPT.
  • 15K video samples from ShareGemini and 15K from Video-ChatGPT, respectively.
  • Our proposed scheme: 10K samples each from ShareGemini and Video-ChatGPT, plus 10K samples synthesized from text data (5K from LongAlpaca and 5K from LongQLora, respectively).
  • Same video samples as above (20K in total), plus 10K samples of corresponding pure text data.
[…]
Notably, replacing the synthetic data with the original pure text counterpart achieves an overall inferior performance. We hypothesize that this is due to the inherent domain gap between vision and text. Thus, to simulate the structure of video frame sequences, transcribing long text into images is necessary.
Figure 7: Scaling performance with different data volumes and types on general video benchmarks. We consider four settings, i.e. fine-tuning with video data only, our hybrid data (video data mixed with our synthetic samples with a ratio of 2:1:212:12 : 1), TOPA data, and T3 data. By default, video data (the caption dataset and the instruction dataset) are mixed with a 1:$1:11$ : 1 sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference. Purple-shaded area indicates gains of our methods over the video data baseline.
Can synthetic data help models understand longer videos? Interestingly, in the training stage, we only utilize synthetic samples of long multimodal context instead of authentic long video samples. However, on the long video benchmark set, our proposed method still achieves a score that is 1.2 points higher than the full data training (as shown in the 4th row compared to 200K full data in Tab. 2). This result suggests that fine-tuning with a long multimodal context can enhance the comprehension of longer videos. In the following section, we will present additional results and discussions to illustrate this point further.
[…]
Baseline 0 24 40.1 40.0 44.5 41.6
30 24 44.7 39.7 45.4 43.3
60 24 46.2 42.7 46.2 45.1
100 24 46.7 44.1 45.3 45.3
Our Method 30 24 45.6(+0.9) 48.7(+9.0) 51.4(+6.0) 48.5(+5.2)
60 24 46.2 51.2(+8.5) 53.2(+7.0) 50.2(+5.1)
100 24 48.7(+2.0) 50.1(+6.0) 57.0(+11.7) 51.9(+6.6)
Table 3: Long video understanding performance with different data volumes and types. The Baseline method adopts only video data, while Our Method utilizes a mix of video data and synthetic samples attained from our method. In the Baseline group, 0 sample indicates zero-shot inference. The performance gains are calculated relative to the video data fine-tuning baseline trained with the same number of samples.
5.3 Examination of Key Properties
In this section, we further examine key properties of our proposed method, including general effectiveness, scaling performance, and effectiveness when adopted in long video understanding scenarios.
5.3.1 General Effectiveness and Scaling Performance
We further verify the proposed scheme’s effectiveness by evaluating our methods on another image-LLM of larger parameter size, i.e. MiniCPM-8B, across different benchmarks. Specifically, through scaling up with different volumes and types of data, we compare our methods against a pure video data baseline, as well as other relevant methods, including TOPA and T3. Both TOPA and T3 first translate vision information into text, such as captions and relations between objects. Then, synthetic text QAs are constructed to simulate video reasoning samples, aiming to transfer temporal reasoning capabilities from text to video. Note that since the original data format of TOPA is largely different from the current paradigm, we design a template to adapt them to the instruction data format (More details are available in Appendix A.2). The results are summarized in Fig. 7.
General effectiveness. When using the same amount of training samples, our methods almost always outperform other methods by a clear margin in all the evaluated benchmarks. Specifically, when using 30K samples, our method achieves an overall accuracy of 52.7, surpassing the baseline by 3.9 points. Notably, it even outperforms the baseline trained with 100K samples by 1.7 points. Similarly, on the MVBench benchmark, with 100K samples, our method attains a 4.3-point absolute gain over the baseline. Overall, the superior performance on different benchmarks showcases the general effectiveness of our proposed method.
Scaling performance. A clear issue with other methods is that they are more prone to performance saturation when scaling up the data budget. For instance, the baseline method uses 60K samples to improve the Video-MME benchmark by 3.1 points, while using 100K samples only achieves another 0.9 points of absolute gains. In contrast, our proposed scheme shows more stable and consistent improvements when scaling up the data volumes compared with other methods. This suggests that ensuring a more diverse distribution of instruction in video training is very important, or otherwise, the learning efficiency may downgrade. To avoid such circumstances, the proposed data augmentation method can be an effective approach.
Discussion: Can we scale up only with synthetic textual samples? An intriguing and highly relevant question is whether we can scale up the synthetic samples without using any real video samples. Since the text is more compact and less redundant than a whole video, training in this way is more economical. Unfortunately, the empirical results show that this is probably unfeasible. As shown in Fig. 7, scaling with synthetic text data (TOPA and T3) shows undesirable characteristics, i.e. this approach can easily reach the saturation point or even slightly downgrade. Other critical issues include the modality gap and special processing of videos in various domains (such as egocentric videos and movies). Besides, since text suffers inevitably from information loss when translated from videos, text data might be better used as a supplement to videos, which helps inject temporal reasoning language prior into the LLM backbones (similar to TOPA or T3) or mixing with video data as a regularization method (as our method does).
Figure 8: Long video understanding performance evaluated with a different number of frames.Video context” denotes the max number of video frames used in training, i.e. 24 frames; “LLM context” denotes the context window of the LLM backbone. Purple-shaded area indicates performance gains of our proposed method over the video data baseline.
5.3.2 Long Video Understanding Performance
We adopt tailored benchmarks to evaluate long video understanding capabilities, including LongVideoBench [49], MLVU-M [50], and the long video set of Video-MME, and report the performance on evaluation sets for the former two benchmarks. Our study focuses on two aspects: (1) performance improvement in terms of long video understanding compared with the video fine-tuning baseline and (2) frame number (multimodal context) generalization ability in the inference stage.
Performance improvement w.r.t. training sample size. As shown in Tab. 3, we observe consistent improvements in long video understanding compared with the video fine-tuning baseline despite an absence of any long video training data. Notably, when using the same number of 100K training samples, our hybrid data scheme improves by 6.6 points over the baseline. We hypothesize this is because the reasoning patterns within the long text can also transfer to video understanding, which exempts the need to curate specific text samples translated from vision information [27, 28].
Performance change w.r.t. frame number. We also investigate if our approach expands the context window. Larger context windows are usually beneficial since video-LLMs derived from large-context LLM backbones often accompany frame number generalization and benefit from inputting more video frames [49, 26]. However, we do not observe this trend. As shown in Fig. 8, the model fine-tuned with long synthetic samples still follows a similar pattern, that is, when performing inference beyond the video context of the training stage, the performance stays relatively stable and does not benefit from more frame input. And when the frame number exceeds the LLM context, the performance plunges to a low level. Thus, promising directions for further improvements may include continued pre-training to expand the LLM context window.
6 Conclusion
In this paper, we propose Sparrow, a data-efficient training scheme for video-LLMs, which enables training with fewer samples and achieving better video understanding performance. This method derives from our empirical findings that the low learning efficiency in data scaling may be ascribed to a limited instruction diversity in the training corpus. Thus, we design an economical data augmentation method that synthesizes video-like samples rich in instruction diversity. Comprehensive experiments demonstrate the general effectiveness and key properties of our proposed method. We hope this paper’s findings can spark more explorations of efficient training and high-quality video training corpora.
Appendix A More Implementation Details
A.1 Answer Judging
We notice that MiniCPM-8B [1] often fails to follow instructions properly when we explicitly ask the model to “Answer with the option’s letter from the given choices directly”, making simple exact matching inaccurate. Specifically, the model often prepends or appends additional text other than the option letters, e.g.Answer: B. Pink.”, or gives additional explanations apart from the answer.
To cope with these issues, we adopt a combination of exact matching and LLM matching for assessment. Specifically, we strip the prefixes such as “Answer:” from the prediction and try to use regular expression matching to find the option letter. When the exact matching scheme fails, we use an LLM (Llama-3.1-8B-Instruct [51]) to find an option closest to the model prediction. When the LLM matching fails, a placeholder outside of the available options (such as “Z”) is returned to denote a wrong answer. Our judging prompt for the LLM is modified from VLMEvalKit [52], as shown in Tab. 4.
A.2 Reproduction Details of Baseline Methods
Due to an inconsistent formulation between the method TOPA [27] and our proposed method, we adapt the implementation for a fair comparison. The original sample comprises a global caption for the whole video and frame-specific information. The frame-related information contains a frame-level caption and some descriptions of key objects in the frame. Thus, we design a prompt template to fit the original textual samples into the unified training format.
A real case of formatting the sample with the devised template is shown in Tab. 5.
System message
You are an AI assistant who will help me to match an answer with several options of a single-choice question.
Prompt
You are provided with a question, several options, and an answer, and you need to find which option is most similar to the answer.
If the meaning of all options is significantly different from the answer, output Z. You should directly output a single uppercase character, such as A, B, C, D (if they are valid options) and Z, and nothing else. Here are two examples.
Example 1:
Question: What is the main object in the image?
Options: A. teddy bear.
B. rabbit.
C. cat.
D. dog.
Answer: a cute teddy bear
Output: A
Example 2:
Question: What is the main object in the image?
Options: A. teddy bear.
B. rabbit.
C. cat.
D. dog.
Answer: Spider
Output: Z
Now here are the question, options, and the answer, you should match and give me the option letter:
Question: {Question}
Options: {Options}
Answer: {Model Answer}
Output:
Table 4: Template for prompting LLM to perform option matching. {Question} is the specific question of a benchmark sample, and {Options} are corresponding choices of the question. {Model Answer} is the raw prediction of MLLMs.
User:
You will be provided with some information about a video, including a global caption for the whole video, global captions for each video frame, and descriptions of key objects in each frame. Answer the questions using the information below.
Video information:
Caption: This video shows the carpentry process. At first, the person sits on a workbench and measures the length of the plank. Then, he uses a saw to cut the wooden plank into multiple pieces. Then, he uses a hammer to nail two pieces of wood together. Finally, he takes a break and drinks water, and leaves the workshop.
Frame information:
Frame 1:
Caption: the person sits on a workbench and holds a hammer in his right hand.
the person: the person is sitting on a workbench in the center of the frame. He is looking at a wooden plank in front of him.
hammer: He has a hammer in his right hand.
workbench: He is sitting on a workbench in the center of the frame.
Frame 2:
Caption: the person takes a chisel from a red toolbox and starts chiseling the plank.
chisel: There are a chisel and a hammer on the workbench in the center of the frame.
the person: the person is holding a chisel in the center of the frame.
red toolbox: There is a red toolbox in the right corner of the frame.
workbench: He is sitting on a workbench in the center of the frame.
Frame 3:
Caption: the person measures the length of the plank with a measuring tape.
the person: the person carries a pencil and a ruler in his right hand.
pencil: There is a pencil in the right hand of the person.
measuring tape: the person is holding a measuring tape in his right hand.
wooden plank: The plank is above the workbench.
Frame 8:
Caption: the person puts his hammer and chisel into the toolbox and leaves the workshop.
toolbox: the person is putting a hammer and a chisel into a toolbox in the center of the frame.
the person: the person is walking toward the door in the center of the frame.
hammer: There are a hammer and a chisel in the right hand of the person.
workshop: the person is leaving the workshop.
Question: Why does the person use a measuring tape in the third frame?
A. To hammer the plank
B. To cut the plank
C. To measure the length of the plank
D. To handle the saw
E. To fix the length of the wooden plank
Answer with the option’s letter from the given choices directly.
Assistant:
C
Table 5: An example of using the template for structuring textual samples into the training format. For simplicity, we only show part of the frame-level information.